Find the centre of curvature of the curve
x=3t y=t^2-6 at (x,y)
1
Expert's answer
2020-11-25T17:51:24-0500

Here "9(y+6)=x^2." Hence "y=\frac{x^2}{9}-6." Hence "y_1=\frac{2x}{9}, y_2=\frac{2}{9}."

Hence center of curvature is ("x-\frac{y_1[1+y_1^2]}{y_2}, y+\frac{1+y_1^2}{y_2})" . Hence required coordinate is

("x-\frac{2x/9}{2/9}(1+\frac{4x^2}{81})", "y+\frac{1+\frac{4x^2}{81}}{2/9}" )= ("-4x^3/81, y+9/2+\frac{2x^2}{9})"=("-4x^3\/81, 3y+16.5" ) .